# for incompressible flows

### v1

@@ 6. An integral functional measuring the logarithmic regularity of the flow
$\frac{d}{dt}\log|\nabla\Phi|$ $\leq\frac{|\nabla\dot{\Phi}|}{|\nabla\Phi|}=\frac{|\nabla(u(\Phi))|}{|\nabla\Phi|}=\frac{|\nabla u|(\Phi)\,|\nabla\Phi|}{|\nabla\Phi|}=|\nabla u|(\Phi)\,$, (6.1)
@@
  • Establish bounds on $\mathcal{G}(\Phi)$ in term of norms of the velocity field. Here (and only here) we use the dynamics, that is, the fact that $\Phi$ is a solution of (ODE). (Subsection 8.1).
  • Use the bounds on $\mathcal{G}(\Phi)$ to show some mild regularity of the flow, more precisely, a quantitative Lusin-Lipschitz estimate. (Subsection 8.2).
  • Show that the Lusin-Lipschitz estimate implies exponential-in-time lower bounds on the mixing scale. (Subsection 8.3).
@@ 8.1. Quantitative estimates for the functional $\mathcal{G}(\Phi)$
$\frac{d}{dt}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,r)}$ $\log\left(1+\frac{|\Phi(t,x)-\Phi(t,y)|}{r}\right)\,dy\leq\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,r)}\frac{|u(t,\Phi(t,x))-u(t,\Phi(t,y))|}{|\Phi(t,x)-\Phi(t,y)|}\,,dy$
@@
$\frac{d}{dt}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,r)}$ $\log\left(1+\frac{|\Phi(t,x)-\Phi(t,y)|}{r}\right)\,dy\lesssim\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,r)}\log\left(1+\frac{|x-y|}{r}\right)\,dy$
@@
$\mathcal{G}(\Phi)\leq\bar{C}(1+\|Du\|_{L^{1}([0,T];L^{p}(\mathbb{T}^{d}))})\,.$ (8.1)
In an informal sense, estimate (8.1) indicates that the regular Lagrangian flow possesses a logarithmic derivative. Alternative approaches involving logarithmic function spaces have been developed in [MR4263701, MR4377866], in the compressible setting in [MR3862947], and through Littlewood–Paley theory in [MR4747245].
@@ 8.2. Quantitative Lusin-Lipschitz regularity of the flow
Fix $\eta>0$. Using Chebyshev’s inequality applied to
@@
Understand how the Lusin-Lipschitz estimate (8.3) implies compactness of regular Lagrangian flows under suitable bounds, and understand how this can be exploited to prove existence of the regular Lagrangian flow.
@@
The above estimates in the case of the space $\mathbb{R}^{d}$ can be localized. When dealing with regular Lagrangian flows with controlled compression, rather than incompressible ones, the estimates depend on the compressibility constant of the flow.
@@ 8.3. Lower bounds on the mixing scale for uniformly-in-time Sobolev velocity fields
where
$M=1+\sup_{0\leq s\leq t}\|\nabla u(s,\cdot)\|_{L^{p}(\mathbb{T}^{2})}\qquad\text{ and }\qquad\beta=\beta(2,p,\eta,T)\,.$
@@
We now recall Vitali covering lemma: Let $\mathcal{F}$ be a family of balls with uniformly bounded radius contained in the torus. Then, there exists a countable, disjoint subfamily $\mathcal{F}^{\prime}$ such that
@@
Exploiting Vitali covering lemma and (8.7) we deduce that
@@ 9. Optimality of the estimates and properties of the regular Lagrangian flow
The preceding example further implies, via the interpolation estimate (4.4), that for smooth velocity fields and initial data, the positive Sobolev norms of the solution may increase exponentially in time. By suitably patching together such solutions and rescaling them to make the exponential growth arbitrarily fast, in [MR3933614] the following statement on loss of regularity for (PDE) has been shown:
@@
The above results, which stand in contrast to the DiPerna–Lions–Ambrosio theory [MR1022305, MR2096794], suggest the following general principle. For velocity fields in $W^{1,p}$, well-posedness of both (PDE) and (ODE) holds all the way down to $p=1$, thereby extending the classical theory valid in the Lipschitz case $p=\infty$. However, the geometric and regularity properties of the flow and of the solution persist only at the level $p=\infty$, that is, in the Lipschitz regime or nearby it.

### v2

@@ 6. An integral functional measuring the logarithmic regularity of the flow
$\frac{d}{dt}\log|\nabla\Phi|$ $\leq\frac{|\nabla\dot{\Phi}|}{|\nabla\Phi|}=\frac{|\nabla(u(\Phi))|}{|\nabla\Phi|}\leq\frac{|\nabla u|(\Phi)\,|\nabla\Phi|}{|\nabla\Phi|}=|\nabla u|(\Phi)\,$, (6.1)
@@
  • Establish bounds on $\mathcal{G}(\Phi)$ in term of norms of the velocity field. Here (and only here) we use the dynamics, that is, the fact that $\Phi$ is a solution of (ODE) (see Subsection 8.1).
  • Use the bounds on $\mathcal{G}(\Phi)$ to show some mild regularity of the flow, more precisely, a quantitative Lusin-Lipschitz estimate (see Subsection 8.2).
  • Show that the Lusin-Lipschitz estimate implies exponential-in-time lower bounds on the mixing scale (see Subsection 8.3).
@@ 7. Some tools from harmonic analysis
The space $L^{1,\infty}(\mathbb{T}^{d})$ is defined as the space of all measurable functions such that the quantity defined in (7.3) is finite.
@@ 8.1. Quantitative estimates for the functional $\mathcal{G}(\Phi)$
$\frac{d}{dt}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,r)}$ $\log\left(1+\frac{|\Phi(t,x)-\Phi(t,y)|}{r}\right)\,dy\leq\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{B(x,r)}\frac{|u(t,\Phi(t,x))-u(t,\Phi(t,y))|}{|\Phi(t,x)-\Phi(t,y)|}\,dy$
@@
[…]
@@
$\mathcal{G}(\Phi)\leq\bar{C}\big(1+\|Du\|_{L^{1}([0,T];L^{p}(\mathbb{T}^{d}))}\big)\,.$ (8.1)
In an informal sense, estimate (8.1) indicates that the regular Lagrangian flow possesses a logarithmic derivative. Alternative approaches involving logarithmic function spaces have been developed in [MR4263701, MR4377866], removing the divergence-free condition in [MR3862947], and through Littlewood–Paley theory in [MR4747245].
@@ 8.2. Quantitative Lusin-Lipschitz regularity of the flow
We now show how a mild regularity property of the regular Lagrangian flow follows from the bound on $\mathcal{G}(\Phi)$. Fix $\eta>0$. Using Chebyshev’s inequality applied to
@@
Understand how the Lusin-Lipschitz estimate (8.3) implies strong compactness of regular Lagrangian flows under suitable bounds, and understand how this can be exploited to prove existence of the regular Lagrangian flow.
@@
The above estimates in the case of the space $\mathbb{R}^{d}$ can be localized. Moreover, they can also be extended to some cases when the divergence-free condition is dropped.
@@ 8.3. Lower bounds on the mixing scale for uniformly-in-time Sobolev velocity fields
for some constant
$\beta=\beta(2,p,\eta,T)$
and where
$M=1+\sup_{0\leq s\leq t}\|\nabla u(s,\cdot)\|_{L^{p}(\mathbb{T}^{2})}\,.$
@@
We now recall Vitali’s covering lemma: Let $\mathcal{F}$ be a family of balls with uniformly bounded radius contained in the torus. Then, there exists a countable, disjoint subfamily $\mathcal{F}^{\prime}$ such that
@@
Exploiting Vitali’s covering lemma and (8.7) we deduce that
@@ 9. Optimality of the estimates and properties of the regular Lagrangian flow
The preceding example further implies, via the interpolation estimate (4.4), that for smooth velocity fields and initial data, the positive Sobolev norms of the solution may increase exponentially in time. By suitably patching together such solutions and rescaling them to make the exponential growth arbitrarily fast, in [MR3933614] the following statement on loss of regularity for (CE) has been shown:
@@
The above results, which stand in contrast to the DiPerna–Lions–Ambrosio theory [MR1022305, MR2096794], suggest the following general principle. For velocity fields in $W^{1,p}$, well-posedness of both (CE) and (ODE) holds all the way down to $p=1$, thereby extending the classical theory valid in the Lipschitz case $p=\infty$. However, the geometric and regularity properties of the flow and of the solution persist only at the level $p=\infty$, that is, in the Lipschitz regime or nearby it.